Assume a semi-annual coupon bond matures in 3 years, has a face value of $1,000, a current market price of $989, and a 5 percent coupon. Which one of the following statements is correct concerning this bond?

Multiple Choice

  • The current coupon rate is greater than 5 percent.

  • The bond is a money market instrument.

  • The bond will pay less annual interest now than when it was originally issued.

  • The current yield exceeds the coupon rate.

  • The bond will pay semi-annual payments of $50 each.

Next Visit question map

Question9of10Total9 of 10

Prev
Answer #1

Face value of the bond = $ 1000

Coupon Rate =5% (Paid semi annually)

Interest Payments = 1000*5% / 2 =50 /2 = $ 25 each semi annually.

Current Market Price = 989

Current Yield(Semi annual) = 25/989 =2.53%,

Annual Yield Rate = 2.53% *2 = 5.06%.

From the above calculations, we can conclude which of the given comments are correct.

Current Coupon Rate is higher than 5% - No

The bond is a money market instrument. - Yes (As it is actively traded).

The bond will pay less annual interest now than when it was originally issued. - No (A bond is a fixed income instrument, annual interest will not change).

The current yield exceeds the coupon rate - Yes (Current Yield is 5.06% whereas coupon rate is 5%).

The bond will pay semi-annual payments of $50 each. - Semi Annual Payment of $ 25 Each